1/3

Дмитрий Николаевич Садовников,  Дмитрий Николаевич Садовников,
Volga, Volga (Stenka Rasin)  Sten’ka Razin
tradukita de N. N. 20  tradukita de Tr. Jacob Robbins

Bag om øen, på den brede underskønne Volgaflod Stenjka Rasins stolte både ud på nye togter stod.

1. From beyond the wooded island To the river wide and free Proudly sailed the arrow-breasted Ships of Cossack yeomanry.

På den første høvding Stenjka sidder selv i bådens stavn, han på kamp og strid ej tænker, har prinsessen i sin favn.

2. On the first is Stenka Razin With a princess at his side, |: Drunken, holds a marriage revel With his beautiful young bride. :|

Men hans mænd er Donkosakker, de er vrede, for de ved, han har glemt dem for en kvinde, for en nat i kærlighed.

3. But behind them rose a whisper, “He has left his sword to woo; |: One short night, and Stenka Razin Has become a woman too!” :|

Stenjka Rasin hører latter, blodet koger i hans bryst Over Volgas brede strømme højt nu lyder Stenjkas røst.

4. Stenka Razin hears the jeering Of his discontented band, |: And the lovely Persian princess He has circled with his hand. :|

....................................... ....................................... ....................................... .......................................

5. His black brows have come together As the waves of anger rise, |: And the blood comes rushing swiftly To his piercing, jet-black eyes. :|

…  …

2/3

....................................... ....................................... ....................................... .......................................

6. “I will give you all you ask for, Life and heart, and head and hand,” |: Echo rolls the pealing thunder Of his voice across the land. :|

....................................... ....................................... ....................................... .......................................

7. “Volga, Volga, mother Volga, Deep and wide beneath the sun, |: You have never seen a present From the Cossack of the Don. :|

....................................... ....................................... ....................................... .......................................

8. And that peace might rule as always All my free-born men and brave, |: Volga, Volga, mother Volga, Volga, make this girl a grave.” :|

“For at ingen splid skal være mellem frie mænd,- tag så Moder Volga, denne kvinde!” og hans arm om hende lå.

9. With a sudden, mighty movement, Razin lifts the beauty high, |: And he casts her where the waters Of the Volga move and sigh. :|

Og han ta’r med stærke arme da sin brud og kaster ned /: i de dybe, mørke bølger, sine nætters salighed. :/ 10. Now a silence like the grave sinks To all those who stand to see, |: And the battle-hardened Cossacks Sink to weep on bended knee. :|

“Men hvad fanden, hvorfor hænger I med næbbet? Bliv kun ved! Dans og syng de røversange, syng for hendes sjælefred!”

11. “Dance, you fool, and men, make merry! What has got into your eyes? |: Let us thunder out a chanty Of a place where beauty lies.” :|

…  …

3/3

Frem fra øen - ud på dybet stævner Stenjka Rasin ud /: med den stærke, stolte både, sejler fra sin døde brud. :/ 12. From beyond the wooded island To the river wide and free, |: Proudly sail the arrow-breasted Ships of Cossack yeomanry. :|

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Danan de N. N. 20.

Arg-478-1037 (2010-03-08 12:16:20)

Aperinte en 1947 en la kant-libro “DLEA-Kantaro” eldonita de Dana Laborista Esperanto-Asocio (DLEA), n-ro 100 kaj en la kantlibro “Store Esperanto-Sangbog”, SEFO 1939, n-ro 31. Vidu la retejon esperantonia. dk sub http: // www. esperantonia. dk/ volga. htm .

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Anglan de Tr. Jacob Robbins en 1921.

Arg-478-1039 (2010-03-09 17:06:16)

Tiu ĉi angligo troviĝas en http: // ingeb. org/ songs/ stenkara. html .